# Distributed Hybrid Sketching for $\ell_{2}$-Embeddings

### v1

@@ I Introduction
Randomization in numerical linear algebra and data science has been a key tool for dimensionality reduction of large matrices and datasets over the past 25 years [2, 3, 4, 5]. This is an interdisciplinary field which is abbreviated to “RandNLA”. One of the main tools in the field is “sketching”, a term that refers to random linear dimension reduction maps. This method provides an effective and computationally efficient randomized approach to addressing problems like matrix factorization [6], eigenvalue computation [7], $k$-means [8, 9], or solving linear systems [10, 11], which are prohibitive in high dimensions and expensive with deterministic methods. The core idea behind sketching is to leverage randomization to create structured and well-conditioned matrices that preserve important properties of the original matrices after compressing them. By applying known algorithms to the sketched matrices, one obtains an approximate solution that is statistically close to the true solution of the problem. A sufficient condition which guarantees that through sketching we obtain a high quality solution to overdetermined systems is the $\ell_{2}$-subspace embedding ($\ell_{2}$-s.e.) property, which says that with high probability the geometry of the system’s basis is preserved after performing the sketching procedure. A widely used class of sketching matrices satisfying this property, are “oblivious subspace embeddings” (OSEs); which are dense randomized linear maps that project vectors to a lower-dimensional space.
More recently, with the increase in daily data generation which is prevalent in many machine learning and statistical inference models, resorting to distributed systems for storage and computations is a necessity [12, 13]. Furthermore, in many applications the data is generated locally and should not be shared with anyone within the distributed network before encrypting it, as privacy and security concerns may arise, e.g. hospitals in a health care network; each of which are comprised of local servers and are collectively operated by a health organization. A prime example with such a framework which has been extensively studied is that of federated learning (FL) [14, 15, 16, 17]. To this end, distributed sketching has been studied in several settings [18, 19, 20, 21, 22], in which a central aggregating server administers a centralized distributed computing network comprised of $k$ servers with whom the coordinator communicates part of the data, and each of them locally perform a sketch or computation that is sent back. The coordinator then decides on a final sketch of the entire dataset or a solution to an optimization problem. We illustrate such a setting in Figure 1, which is what we study in this paper.
@@
In this article, we focus on the task of distributed $\ell_{2}$-subspace embedding through sketching, as well as its composition with hybrid sketching for enhanced embedding results of distributed data matrices. The main contributions of our work are that we show how we can obtain a global $\ell_{2}$-s.e. through decentralized sketching, and provide the first theoretical analysis of hybrid sketching. We concretely show that in certain cases hybrid sketching is in fact the appropriate sketching technique to consider. Moreover, our proposed framework ensures that no information about the data can be revealed when potential eavesdroppers are present, nor to the central coordinator. That is, we ensure security of the local data, which is desirable in FL. We also discuss the implication our sketching technique has in distributed gradient descent (GD), arguably the most common application of distributed sketching.
@@ I-A Contributions
The contributions of this paper are the following. $(\mathrm{i})$ First, we present a general framework for a global $\ell_{2}$-s.e. through local sketching (Thm. 2), which is more general than related works [23]. $(\mathrm{ii})$ We then conduct a rigorous analysis of hybrid sketching (Thm. 3), which was only considered empirically in [20]. $(\mathrm{iii})$ To this extent, we show that by sequentially applying two distinct $\ell_{2}$-s.e. sketches; we obtain a composite sketching matrix, which applies beyond our distributive hybrid sketching framework. $(\mathrm{iv})$ Finally, what is arguably the most interesting contribution of this work in terms of theoretical results, is that by a judicious choice of locally sketching through Subsampled Randomized Hadamard Transforms (SRHTs) and by then applying a Rademacher sketching matrix on the aggregated sketch globally through, it is possible to benefit in terms of operation and/or reduced dimension in comparison to other prevalent sketching matrices (Sec. V-B). This is a consequence of the fact that we use fast OSEs locally, and an OSE with better embedding properties globally. This approach is also numerically superior in terms of the $\ell_{2}$-s.e. error.
The motivation of our proposed hybrid scheme, was to concurrently satisfy the following desiderata, to further accelerate computations that rely on large distributed datasets: $(\mathrm{a})$ security in distributed data aggregation, $(\mathrm{b})$ local compression of data aggregation; while maintaining the geometric properties of the global dataset, $(\mathrm{c})$ perform the compression efficiently and distributively, $(\mathrm{d})$ reduce the communication load required to deliver the local data, and $(\mathrm{e})$ reduce the space complexity required by the central server. Although for our theory we focus on least squares; a fundamental primitive tool used for solving a variety of optimization and estimation problems, our approach is applicable in other downstream applications.
@@ I-B Related Work
The closest works to what we present, are distinct sketching techniques from [23, 21, 20]. In a recent monograph [24], the work of [21] is the only one mentioned in which local sketching is performed. In [21], which coined the term “block-SRHT [^1] We note that a technique in a different context was also called “block-SRHT[19], which we do not explicitly discuss in this article., the authors assume that a partition of the global dataset is sent to each of the nodes, who then apply a SRHT to their allocated partition; along with an additional global permutation and a local signature matrix, before sending back the sketch to the coordinator who sums the sketches.
@@
The work of [23] develops a technique that the authors term “Sparse-SRHT” for linear regression, which resembles the algorithm we present in Sec. IV. Our algorithm can be viewed as a generalization of theirs, as they only consider local sketching performed through a regular SRHT, and do not consider the distributed setting depicted in Figure 1. Furthermore, they assume that the data is centrally administered and do not leverage the fact that the computation, which is true also in their case, can be divided and performed in parallel by $k$ server nodes. We do however use a result of theirs to simplify our comparisons in hybrid sketching, primarily to avoid the use of an oversampling parameter, which is an artifact of the techniques we are using for our distributed algorithm. One difference between our approach and theirs, is that they also consider an additional permutation matrix that is performed on the global dataset $\mathbf{A}$, which requires either further communication between the servers or that the dataset is aggregated before local sketching takes place. As they explain, this does not affect the theoretical analysis, though we observed that experimentally such a permutation improves the flattening of the transformed leverage scores.
@@
Other related work include distributed sketching techniques in which different sketches are performed on the global dataset; to solve varying sketched versions of the global system of linear equations, which are then averaged or aggregated to determine a good approximate solution for linear regression [18, 20]. The main drawback here compared to our approach, is that each sketch is performed on the global dataset; and therefore it is not suitable for systems in which local nodes may wish to not share their information with other nodes.
@@
Organization: The paper is organized as follows. In Sec. II, we recall the definition of an $\ell_{2}$-s.e., and give an outline of our proposed scheme which implements such an embedding distributively without aggregating the data a priori. In Sec. III we setup the necessary notation and review background from RandNLA which we will need for the main portion of the paper. We then move to the main body, where in Sec. IV we formally present our algorithm for distributed local sketching; and how it obtains a global $\ell_{2}$-s.e. The analyses of our results are presented in Sec. IV-A. In Sec. V we extend the work of Sec. IV to hybrid sketching. Finally, we present numerical experiments in Sec. IV-D and Sec. VI; and concluding remarks in VII.
@@ II Problem Setup and Our Distributed Scheme
One of the most representative applications of RandNLA is the linear least squares approximation problem, in which we seek to approximately solve the overdetermined linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ by solving
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n\times d}$ and $\mathbf{b}_{i}\in\mathbb{R}^{n}$ for all $i$, and $n=N/k$; for which $n>d$. We consider the reduced SVD of $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{\mathbb{N}\times d}$ and $\mathbf{A}$ is full rank. To simplify our presentation, we assume that $k|N$; and $\{\mathbf{A}_{\iota}\}_{\iota=1}^{k}$ are equipotent. A way to approximate (1) in a faster manner, is to instead solve the modified least squares problem
${\hat{\mathbf{x}}}=\arg\min_{\mathbf{x}\in\mathbb{R}^{d}}\Big{\{}L_{\mathbf{S}}(\mathbf{A},\mathbf{b};\mathbf{x})\coloneqq\|\mathbf{S}(\mathbf{A}\mathbf{x}-\mathbf{b})\|_{2}^{2}\Big{\}}$ (3)
@@
A sketching matrix $\mathbf{S}\in\mathbb{R}^{R\times N}$ is a $\ell_{2}$-subspace embedding (or satisfy the $\ell_{2}$-s.e. property) of $\mathbf{A}\in\mathbb{R}^{N\times d}$ with a left orthonormal basis $\mathbf{U}$, and $N\gg d$; $N>R>d$, if for any $\mathbf{y}\in\mathsf{im}(\mathbf{U})$ we have with high probability:
@@
In turn, Definition 1 characterizes the approximation’s error of the solution ${\hat{\mathbf{x}}}$ of (3) as
@@
and $\|\mathbf{A}(\mathbf{x}^{\star}-{\hat{\mathbf{x}}})\|_{2}\leqslant\epsilon\|(\mathbf{I}_{N}-\mathbf{U}\mathbf{U}^{\top})\mathbf{b}\|_{2}$ [35]. Furthermore, desired properties of sketching matrices are that they are zero-mean and normalized, i.e. $\mathbb{E}[\mathbf{S}]=\mathbb{0}_{R\times N}$ and $\mathbb{E}[\mathbf{S}^{\top}\mathbf{S}]=\mathbf{I}_{N}$ [36], which are met by normalized random matrices with i.i.d. standard Gaussian entries.
In terms of $\ell_{2}$-s.e., we show that it is possible to aggregate sketches of local datasets from a distributed network, to obtain a sketch of the collective dataset as a whole (referred to as the “global dataset”). We work with OSEs, a category of $\ell_{2}$-s.e. which suit our objective. Specifically, by performing a random projection on the local datasets $\{\mathbf{A}_{i}\}_{i=1}^{k}$, we show that in the resulting global dataset $\mathbf{A}$ the transformation of each data point is of approximately equal importance; which is quantified by leverage scores. Thereby, this implies that local uniform sampling suffices. Embeddings in such data-distributed settings are useful in a variety of data pre-processing/training and unsupervised learning tasks such as clustering and PCA [37, 38, 39, 40, 41].
To obtain the sketch of the global dataset, the distributed nodes send their local sketches $\{\mathbf{S}_{i}\mathbf{A}_{i}\}_{i=1}^{k}$ to a coordinator who aggregates them, to obtain a global sketch $\widehat{\mathbf{A}}\in\mathbb{R}^{R\times d}$ where $R\ll N$. Depending on the choice of the random projection performed by the nodes, there is a security guarantee on their local information; which prohibits potential eavesdroppers and the coordinator from recovering the data points, which is of increasing importance in distributed machine learning. Ultimately, we obtain a summary of the global dataset in a decentralized manner, without explicitly revealing or aggregating the local data. This approach is depicted pictorially in Figure 1.
@@
In our work, a second sketching matrix $\mathbf{\Phi}$ is applied to the global sketch $\widehat{\mathbf{A}}$, to obtain further compression through hybrid sketching. Depending on the combination of sketching matrices and projections considered for the pairs $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$, hybrid sketching may lead to better performance results than when applying a single sketching matrix of the same final target dimension. Specifically, the ideal combination is to first apply a fast sketching matrix (which in our case is done distributively through local sketching), and then a second one with better properties in terms of dimensionality reduction. This technique leads to further benefits on distributed sketching, and is an improvement on the work of [20]; which only considered uniform sampling for $\mathbf{S}_{\{k\}}$ and did not provide any theoretical justification to their approach. Surprisingly, in numerical experiments there was a clear distinction in terms of the empirical embedding guarantee of hybrid sketching versus standard sketching techniques.
In Table I we summarize the most relevant results regarding OSE sketches, which we will use in Sec. V for comparison to hybrid sketching, and to show when there is an improvement through our proposal. We also include sketching through leverage score sampling for completeness, which is the main tool we use for proving Theorem 2, though we do not consider sparse sketches. For further details on these or other sketching techniques, refer to the monographs [3, 42, 2, 43, 44, 5, 24, 45, 46, 47].
@@
Leverage Scores [53, 35] $\Theta\big{(}(d\log(d/\delta)/(\beta\epsilon^{2})\big{)}$ $\mathcal{O}(Rd)$
Hybrid (Radem. &\&& SRHTs) $\mathcal{O}\big{(}(d+\log(1/\delta))/\epsilon^{2}\big{)}$ $\mathcal{O}\big{(}Nd(\log(n)+N\mu^{2}\rho)\big{)}$
TABLE I: The $1^{st}$ column indicates the sketching technique. The $2^{nd}$ column shows the reduced embedding dimension of the resulting sketch, i.e. $R$ for $\mathbf{S}\in\mathbb{R}^{R\times N}$; required for (5) to be satisfied. The $3^{rd}$ column indicates the multiplication time it takes to apply $\mathbf{S}$ on $\mathbf{A}$, with naive matrix-matrix multiplication. ROS in the $3^{rd}$ row stands for “Randomized Orthonormal Systems”. Parameter $\beta\in(0,1]$ in the $6^{th}$ row, is an misestimation factor that accounts for sampling according to an approximate distribution. The compression factors of the first and second sketch respectively in hybrid sketching, are respectively $\mu,\rho\in(0,1)$.

### v2

@@ I Introduction
Randomization in numerical linear algebra and data science has been a key tool for dimensionality reduction of large matrices and datasets over the past 25 years [2, 3, 4, 5]. This is an interdisciplinary field which is abbreviated to “RandNLA”. One of the main tools in the field is “sketching”, a term that refers to random linear dimension reduction maps. This method provides an effective and computationally efficient randomized approach to addressing problems like matrix factorization [6], eigenvalue computation [7], $k$-means [8, 9], or solving linear systems [10, 11], which are prohibitive in high dimensions and expensive with deterministic methods. The core idea behind sketching is to leverage randomization to create structured and well-conditioned matrices that preserve important properties of the original matrices after compressing them. By applying known algorithms to the sketched matrices, one obtains an approximate solution that is statistically close to the true solution of the problem. A sufficient condition which guarantees that through sketching we obtain a high quality solution to overdetermined systems is the $\ell_{2}$-subspace embedding ($\ell_{2}$-s.e.) property, which says that with high probability the geometry of the system’s basis is preserved after performing the sketching procedure. A widely used class of sketching matrices satisfying this property are “oblivious subspace embeddings” (OSEs) which are dense randomized linear maps that project vectors to a lower-dimensional space.
More recently, with the increase in daily data generation which is prevalent in many machine learning and statistical inference models, resorting to distributed systems for storage and computations is a necessity [12, 13]. In many applications the data is generated locally and should not be shared with anyone within the distributed network before encrypting it, as privacy and security concerns may arise, e.g. hospitals in a health care network each of which are comprised of local servers and are collectively operated by a health organization. A prime example with such a framework which has been extensively studied is that of federated learning (FL) [14, 15, 16, 17]. To this end, distributed sketching has been studied in several settings [18, 19, 20, 21, 22], in which a central aggregating server administers a centralized distributed computing network comprised of $k$ servers with whom the coordinator communicates part of the data, and each of them locally perform a sketch or computation that is sent back. The coordinator then decides on a final sketch of the entire dataset or a solution to an optimization problem. We illustrate such a setting in Fig. 1, which we study.
@@
The main benefits of hybrid sketching are that we can obtain high quality sketches when: ($\mathrm{I}$) the data is either inherently distributed and too large to gather into one place or store at a single server before sketching takes place; ($\mathrm{II}$) the locally generated data is sensitive, and local sketching can provide a level of privacy. Through hybrid sketching we can maximize the information of the local data retained when bandwidth and other network constraints are imposed, so that the resulting sketch can have better approximation guarantees compared to distributed local sketching. Throughout the article, other benefits of hybrid sketching will also be discussed.
In this paper, we focus on the task of distributed $\ell_{2}$-subspace embedding through sketching, as well as its composition with hybrid sketching for enhanced embedding results of distributed data matrices. The main contributions of our work are that we show how we can obtain a global $\ell_{2}$-s.e. through decentralized sketching, and provide the first theoretical analysis of hybrid sketching. We concretely show that in certain cases hybrid sketching is in fact the appropriate sketching technique to consider. Moreover, our proposed framework ensures that no information about the data can be revealed when potential eavesdroppers are present, nor to the central coordinator. That is, we ensure security of the local data, which is desirable in FL. We also discuss the implication our sketching technique has in distributed gradient descent (GD), arguably the most common technique used in machine learning.
@@ I-A Contributions
The contributions of this paper are the following. $(\mathrm{i})$ First, we present a general framework for a global $\ell_{2}$-s.e. through local sketching (Thm. 2), which is more general than related works [23]. $(\mathrm{ii})$ We then conduct a rigorous analysis of hybrid sketching (Thm. 3), which was only considered empirically in [20]. $(\mathrm{iii})$ To this extent, we show that by sequentially applying two distinct $\ell_{2}$-s.e. sketches we obtain a composite sketching matrix, which applies beyond our distributive hybrid sketching framework. $(\mathrm{iv})$ Finally, what is arguably the most interesting contribution of this work in terms of theoretical results, is that by a judicious choice of locally sketching through Subsampled Randomized Hadamard Transforms (SRHTs) and by then applying a Rademacher sketching matrix on the aggregated sketch globally through, it is possible to benefit in terms of multiplication time and/or reduced dimension in comparison to other prevalent sketching matrices (Sec. V-B). This is a consequence of the fact that we use fast OSEs locally, and an OSE with better embedding properties globally. This approach is also numerically superior in terms of the $\ell_{2}$-s.e. error.
The motivation of our proposed hybrid scheme, was to concurrently satisfy the following desiderata, to further accelerate computations that rely on large distributed datasets: $(\mathrm{a})$ security in distributed data aggregation, $(\mathrm{b})$ local compression of data aggregation while maintaining the geometric properties of the global dataset, $(\mathrm{c})$ perform the compression efficiently and distributively, $(\mathrm{d})$ reduce the communication load required to deliver the local data, and $(\mathrm{e})$ reduce the space requirement required by the central server. Although for our theory we focus on least squares, a fundamental primitive tool used for solving a variety of optimization and estimation problems, our approach is applicable in other downstream applications which we numerically validate.
@@ I-B Related Work
The closest works to what we present, are distinct sketching techniques from [23, 21, 20]. In a recent monograph [24], the work of [21] is the only one mentioned in which local sketching is performed. In [21], which coined the term “block-SRHT”, the authors assume that a partition of the global dataset is sent to each of the nodes, who then apply a SRHT to their allocated partition along with an additional global permutation and a local signature matrix, before sending back the sketch to the coordinator who sums the sketches.
@@
The work of [23] develops a technique that the authors term “Sparse-SRHT” for linear regression, which resembles the algorithm we present in Sec. IV. Our algorithm can be viewed as a generalization of theirs, as they only consider local sketching performed through a regular SRHT, and do not consider the distributed setting depicted in Fig. 1. Furthermore, they assume that the data is centrally administered and do not leverage the fact that the computation, which is true also in their case, can be divided and performed in parallel by $k$ server nodes. We do however use a result of theirs to simplify our comparisons in hybrid sketching, primarily to avoid the use of an oversampling parameter, which is an artifact of the techniques we are using for our distributed algorithm. One difference between our approach and theirs, is that they also consider an additional permutation matrix that is performed on the global dataset $\mathbf{A}$, which requires either further communication between the servers or that the dataset is aggregated before local sketching takes place. As they explain, this does not affect the theoretical analysis, though we observed that experimentally such a permutation improves the flattening of the transformed leverage scores.
@@
Other related work include distributed sketching techniques in which different sketches are performed on the global dataset to solve varying sketched versions of the global system of linear equations, which are then averaged or aggregated to determine a good approximate solution for linear regression [18, 20]. The main drawback here compared to our approach is that each sketch is performed on the global dataset, and therefore it is not suitable for systems in which local nodes may wish to not share their information with other nodes.
@@
Organization: The paper is organized as follows. In Sec. II, we recall the definition of an $\ell_{2}$-s.e., and give an outline of our proposed scheme which implements such an embedding distributively without aggregating the data a priori. In Sec. III we setup the necessary notation and review background from RandNLA which we will need for the main portion of the paper. We then move to the main body, where in Sec. IV we formally present our algorithm for distributed local sketching and how it obtains a global $\ell_{2}$-s.e. The analyses of our results are presented in Sec. IV-A. In Sec. V we extend the work of Sec. IV to hybrid sketching. In Sec. VI we discuss the implications of our proposed methods to distributed GD, as well as their security and privacy aspects. Finally, we present numerical experiments in Sec. VII-A and Sec. VII, and concluding remarks in Sec. VIII. All proofs can be found in Appendix A, and in Appendix B we discuss an intuitive justification for our proposed hybrid approach.
@@ II Problem Setup and Our Distributed Scheme
One of the most representative applications of RandNLA is the linear least squares problem, in which we seek to approximately solve the overdetermined linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$:
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n\times d}$ and $\mathbf{b}_{i}\in\mathbb{R}^{n}$ for all $i$, and $n=N/k$ for which $n>d$. We consider the reduced SVD of $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{\mathbb{N}\times d}$ and $\mathbf{A}$ is full rank. To simplify our presentation we assume that $k|N$, and $\{\mathbf{A}_{\iota}\}_{\iota=1}^{k}$ are equipotent. A way to approximate (1) in a faster manner, is to instead solve the modified least squares problem
${\hat{\mathbf{x}}}=\arg\min_{\mathbf{x}\in\mathbb{R}^{d}}\Big\{L_{\mathbf{S}}(\mathbf{A},\mathbf{b};\mathbf{x})\coloneqq\|\mathbf{S}(\mathbf{A}\mathbf{x}-\mathbf{b})\|_{2}^{2}\Big\}\$(3)
@@
A sketching matrix $\mathbf{S}\in\mathbb{R}^{R\times N}$ is a $\ell_{2}$-subspace embedding (or satisfy the $\ell_{2}$-s.e. property) of $\mathbf{A}\in\mathbb{R}^{N\times d}$ with a left orthonormal basis $\mathbf{U}$, and $N\gg d$ s.t. $N>R>d$, if for any $\mathbf{y}\in\mathsf{im}(\mathbf{U})$ we have with high probability:
@@
In turn, Definition 1 characterizes the approximation’s error of the solution $^{\hat{\mathbf{x}}}$ of (3) as
@@
and $\|\mathbf{A}(\mathbf{x}^{\star}-{\hat{\mathbf{x}}})\|_{2}\leqslant\epsilon\|(\mathbf{I}_{N}-\mathbf{U}\mathbf{U}^{\top})\mathbf{b}\|_{2}$ [35]. Furthermore, desired properties of sketching matrices are that they are zero-mean and normalized, i.e. $\mathbb{E}[\mathbf{S}]=\mathbb{0}_{R\times N}$ and $\mathbb{E}[\mathbf{S}^{\top}\mathbf{S}]=\mathbf{I}_{N}$ [36] to guarantee unbiased estimates, which are met by normalized random matrices with i.i.d. standard Gaussian entries. Intuitively, random matrices $\mathbf{S}$ (e.g. Gaussian, Rademacher, SRHT) acts like scramblers, and spread out information across the rows of $\mathbf{S}\mathbf{A}$ uniformly. When $\mathbf{S}$ is applied to $\mathbf{A}$, the information in any low-dimensional subspace of $\mathbb{R}^{n}$ does not collapse nor is significantly distorted.
In terms of $\ell_{2}$-s.e., we show that it is possible to aggregate sketches of local datasets from a distributed network, to obtain a sketch of the collective dataset as a whole (referred to as the “global dataset”). We work with OSEs, a category of $\ell_{2}$-s.e. which suit our objective. Specifically, by performing a random projection on the local datasets $\{\mathbf{A}_{i}\}_{i=1}^{k}$, we show that in the resulting global dataset $\mathbf{A}$ the transformation of each data point is of approximately equal importance, which is quantified by leverage scores. Thereby, this implies that local uniform sampling suffices. Embeddings in such data-distributed settings are useful in a variety of data pre-processing/training and unsupervised learning tasks such as clustering and PCA [37, 38, 39, 40, 41].
To obtain the sketch of the global dataset, the distributed nodes send their local sketches $\{\mathbf{S}_{i}\mathbf{A}_{i}\}_{i=1}^{k}$ to a coordinator who aggregates them, to obtain a global sketch $\widehat{\mathbf{A}}\in\mathbb{R}^{R\times d}$ where $R\ll N$. Depending on the choice of the random projection performed by the nodes, there is a security guarantee on their local information which prohibits potential eavesdroppers and the coordinator from recovering the data points, which is of increasing importance in distributed machine learning. Ultimately, we obtain a summary of the global dataset in a decentralized manner, without explicitly revealing or aggregating the local data. This approach is depicted pictorially in Fig. 1.
@@
In our work, a second sketching matrix $\mathbf{\Phi}$ is applied to the global sketch $\widehat{\mathbf{A}}$, to obtain further compression through hybrid sketching. Depending on the combination of sketching matrices and projections considered for the pairs $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$, hybrid sketching may lead to better performance results than when applying a single sketching matrix of the same final target dimension. Specifically, the ideal combination is to first apply a fast sketching matrix (which in our case is done distributively through local sketching), and then a second one with better properties in terms of dimensionality reduction. This technique leads to further benefits on distributed sketching, and is an improvement on the work of [20] which only considered uniform sampling for $\mathbf{S}_{\{k\}}$ and did not provide any theoretical justification to their approach. Surprisingly, in numerical experiments there was a clear distinction in terms of the empirical embedding guarantee of hybrid sketching versus standard sketching techniques.
In Table I we summarize the most relevant results regarding OSE sketches, which we will use in Sec. V for comparison to hybrid sketching, and to show when there is an improvement through our proposal. We also include sketching through leverage score sampling for completeness, which is the main tool we use for proving Theorem 2, though we do not consider sparse sketches. For further details on these sketching techniques, refer to the monographs [3, 42, 2, 43, 44, 5, 24, 45, 46, 47].
@@
Leverage Scores [35, 22] $\Theta\big(d\log(d/\delta)/(\beta\epsilon^{2})\big)$ $\mathcal{O}(Rd)$
Hybrid (Radem. &\& SRHTs) $\mathcal{O}\big((d+\log(1/\delta))/\epsilon^{2}\big)$ $\mathcal{O}\big(Nd(\log(n)+N\mu^{2}\rho)\big)$
TABLE I: The $1^{st}$ column indicates the sketching technique. The $2^{nd}$ column shows the reduced embedding dimension of the resulting sketch, i.e. $R$ for $\mathbf{S}\in\mathbb{R}^{R\times N}$ required for (5) to be satisfied. The $3^{rd}$ column indicates the multiplication time it takes to apply $\mathbf{S}$ on $\mathbf{A}$, with naive matrix-matrix multiplication. ROS in the $3^{rd}$ row stands for “Randomized Orthonormal Systems”. Parameter $\beta\in(0,1]$ in the $6^{th}$ row, is an misestimation factor that accounts for sampling according to an approximate distribution. The compression factors of the first and second sketch respectively in hybrid sketching, are respectively $\mu,\rho\in(0,1)$.